OpenStudy (anonymous):

Can someone tell me if I got this right PLEASE! 8+4x=10+x The answer I got was 4.5

jimthompson5910 (jim_thompson5910):

4.5 is incorrect, try it again

jimthompson5910 (jim_thompson5910):

If you're completely stuck, then I'll get you started 8+4x=10+x 8+4x-x=10 4x-x=10-8 ... ... ... x = ???

OpenStudy (anonymous):

Is it 2?

jimthompson5910 (jim_thompson5910):

8+4x=10+x 8+4x-x=10 4x-x=10-8 3x = 2 x = 2/3

jimthompson5910 (jim_thompson5910):

so not sure how you got 4.5 and 2, but hopefully that will clear things up

OpenStudy (anonymous):

Thanks so much but how is it 2/3?

jimthompson5910 (jim_thompson5910):

well I showed how it's 2/3 in the steps above, which line doesn't make sense? Please let me know. Thanks

jimthompson5910 (jim_thompson5910):

the basic idea is to isolate x, which is done by moving non x terms to one side and the x terms to the other from there you combine like terms to get 3x = 2 last step is to divide both sides by 3 to isolate x to get x = 2/3

OpenStudy (anonymous):

we need to find x to find it we isolate it 8+4x=10+x so we move 8 on RHS and x on LHS +4x-x=10-8 3x = 2 x = 2/3 hope this helps :)

OpenStudy (anonymous):

x = (2/3)